1. Spasmodic dysphonia (SD) is an idiopathic, chronic, and disabling voice disorder. In a case-control study, 150 cases of SD were identified along with 150 controls. Of those with SD, 16 had previously experienced meningitis. Of the controls, 3 had previously experienced meningitis.

a. Which measure of association/comparison should we use to assess the strength of the relationship between the exposure and disease?

b. Calculate the correct measure and show your work.

c. How do we interpret this effect estimate?
Answer #1

Firstly we will deduce the information into the table:

Cases (SD) Cases(NO SD)
Meningitis 16 3
No Meningitis 134 147

(a) By measuring and comparing the incidence of the outcome of interest in two or more groups categorized by extent of exposure, we can begin to assess whether there is an association between exposure and outcome.

  • What type of study design is this? -> Case Control

(b)

Calculate the appropriate estimate of risk between the exposure and the disease?

16x147/3x134=2352/402=5.9

(c)The odds of developing SD after being exposed to meningitis is 5.9 times more likely than those who did not have meningitis.